OpenStudy (anonymous):

lim x>infinite 1/x^2 S(integral sign) tan^-1 dt

OpenStudy (ash2326):

\[\lim_{x\to \infty} \frac 1 {x^2} \int \tan^{-1} t\ dt\]Is this your question?

OpenStudy (anonymous):

yes sorry Im new to this :-)

OpenStudy (ash2326):

No need to be sorry :)

OpenStudy (ash2326):

The integral doesn't involve x, so it can be taken out. \[\int \tan^{-1} t\ dt\times \lim_{x\to \infty} \frac 1 {x^2} \] tell me what's \[\frac 1 \infty=???\]

OpenStudy (ash2326):

@tjutta ???

OpenStudy (anonymous):

\[\infty\]

OpenStudy (ash2326):

nope \[\frac 1 0= \infty\] so \[\frac 1 \infty=???\]

OpenStudy (anonymous):

0

OpenStudy (ash2326):

good :D \[\int \tan^{-1} t\times 0=???\]

OpenStudy (anonymous):

SIN/COS ? :-)

OpenStudy (ash2326):

if you multiply 0 by tan t then what you'd get?

OpenStudy (anonymous):

0

OpenStudy (ash2326):

so it's 0

OpenStudy (anonymous):

what ? the answer is 0 :-) Ok it makes since if you put it that way but for me whos math retarded its difficult:-)

OpenStudy (ash2326):

You know the integral is a variable of t, if you integrate it and multiply by zero. You'd get zero. So it's not necessary to integrate. It's 0

OpenStudy (ash2326):

I doubt that this is so easy. I think you have missed some portion of the question. Maybe limits on the integral. Could you check the question again?

OpenStudy (anonymous):

its just as you put it so I think this is right but I have one more :-) its \[\int\limits_{}^{} \frac{ (x^3)\ }{(x-1)^2(x+2) }\] Can you help me with that?

OpenStudy (ash2326):

okay, did you try this?

OpenStudy (anonymous):

I dont know what to do :-) should I start by divide ?

OpenStudy (ash2326):

What's the expansion of (x-1)^2?

OpenStudy (anonymous):

(x-1)(x+1) ?

OpenStudy (ash2326):

What would you get if you multiply these?

OpenStudy (anonymous):

ok I only have 2 ?

OpenStudy (ash2326):

\[(x-1)\times (x-1)=???\]

OpenStudy (anonymous):

x^2 - 2x + 1

OpenStudy (ash2326):

Good:) Numerator we have x^2, so add and subtract (-2x+1) Could you rewrite the integral using this?

OpenStudy (anonymous):

\[\int\limits_{}^{} \frac{ x }{-2x+1 } \] Im I on the right way? I know Im not very good at this at all :-)

OpenStudy (ash2326):

I'm sorry, i thought it's x^2 in numerator Sorry. Let me think about this